the father is four times older than his son, if their gap age is 21, what are their age?
Answer by ramkikk66(644)

Let son's age be x
Father's age = 4x
Gap in ages = 4x - x = 3x
So 3x = 21
x = 21/3 = 7
Son's age = 7 and father's age = 7*4 = 28.
:)